Need help with #9 I know the answer is C but I’m not getting that answer when I try and solve it
s. weighs 20 N and is held wire? up by a wire as shown below. What is the tension in the A) 60N B) 50 N 45 N 25 N 20 N sin53 Sns3 530 Suppose you pour water into a container until is reaches a depth of 40 cm. Next, you pour in a 30 cm thickness oil so that it floats on the water. If the density of the oil is 800 kg/m, what is the pressure at the bottom of the container due to just the oil and water 9 A) 2,400 N/m2 仁-+ R.Ygxn.usert faxa xhoi i 4,000 N/m2 C 6,400 N/m2 D) 7,000 N/m2 E) 28,140 N/m2 condensed oir density He A metal necklace worn by a baseball player has a mass of 200 grams. When it is placed in water, it sinks and displaces 50 grams of water. The density of this metal necklace is: A) 1.33 g/cm B) 2 g/cm 509 HLO has vol. = 50 cm3 3 g/cm 4 g/cm 5 g/cm volume of Neck 5 0 A 60 kg student sits on a 4 kg chair, Each leg of the chair makes contact with the floor in a circle that has a radius of 1.5 cm. ICalculate the pressure exerted on the floor by each leg of the chair. 10110 452 x105 N/㎡ 3.39 x105 Nm2 2.26 x10 N/m2 13 x10 N/m B) 千兀(45)
Answer #1

9. Pressure at the bottom due to oil and water = P

now height of oil column, ho = 0.3 m

height of water column, hw = 0.4 m

density of water = 1000 kg/m^3 = rhow

density of oil = 800 kg/m^3 = rhoo

hence

P = (rhow*hw + rhoo*ho)g

P = 9.81(1000*0.4 + 800*0.3)

P = 6278.4 Pa

hence

option C 6400 N/m^2